Question 3: A random variable X has a Bernoulli distribution with parameter θ є (0,1) if X {0,1} and P(X-1)-θ. Suppose that we have nd random variables y, x, following a Bernoulli(0) distribution and observed values y1,... . Jn a) Show that EIX) θ and Var[X] θ(1-0). b) Let θ = ỹ = (yit . .-+ yn)/n. Show that θ is unbiased for θ and compute its variance. c) Let θ-(yit . . . +yn + 1)/(n + 2) (this estimator comes from something called Laplaces rule d) Compute the risk (bias squared plus variance) of θ and Comment on the quality of the two Note: Bias is the difference between the expected value of an estimator and the parameter it is of succession). Show that is based but that it has less variance that θ estimators for θ. estimating
Answer #1

a)

E[X] = 0 * P(X = 0) + 1 * P(X = 1) = 1 * \theta = \theta

E[X2] = 02 * P(X = 0) + 12 * P(X = 1) = 1 * \theta = \theta

Var[X] = E[X2] - E[X]2  = \theta - \theta^2 = \theta(1 - \theta)

b)

E[\hat{\theta}] = E[\bar{y}] = E[y] = \theta (By central limit theorem, E[\bar{y}] = E[y] )

Thus, \hat{\theta} is unbiased estimator of \theta

Var[\hat{\theta}] = Var[\bar{y}] = Var[y]/n = \theta(1- \theta)/n

c)

E[\tilde{\theta}] = E[(y_1+ ..+y_n+1)/(n+2)] = E[(y_1+ ..+y_n)+1]/(n+2)

= E[n((y_1+ ..+y_n)/n)+1]/(n+2) = E[n\bar{y}+1]/(n+2) = (nE[\bar{y}] + 1)/n+2

E[\tilde{\theta}] = (nE[y] + 1)/n+2 = (n \theta + 1)/(n+2)

As, E[\tilde{\theta}] \ne \theta , \tilde{\theta} is biased estimator of \theta.

Var[\tilde{\theta}] = Var[(y_1+ ..+y_n+1)/(n+2)] = Var[(y_1+ ..+y_n)+1]/(n+2)^2

= Var[n((y_1+ ..+y_n)/n)+1]/(n+2)^2 = Var[n\bar{y}+1]/(n+2)^2

= (n^2Var[\bar{y}])/(n+2)^2 = (n^2 Var[y]/n) / (n+2)^2 = n \theta (1- \theta)/(n+2)^2

Var[\tilde{\theta}] = n^2 \theta (1- \theta)/n(n+2)^2 = (n / n+2)^2 Var[\hat{\theta}]

For n > 1, (n / n+2)2 < 1

thus,

Var[\tilde{\theta}] < Var[\hat{\theta}]

d)

Bias for \hat{\theta} is,

b = E[\hat{\theta}] - \theta = \theta - \theta = 0

Risk = b2 + Var(\hat{\theta}) = (1-0)/11

Bias for \tilde{\theta} is,

b = E[\tilde{\theta}] - \theta =(n \theta + 1)/(n+2) - \theta = ( 1 - 2\theta)/(n+2)

Risk = b2 + Var(\tilde{\theta}) = ( 1 - 2\theta)^2/(n+2)^2 + n \theta (1- \theta)/(n+2)^2 = (( 1 - 2\theta)^2 +n \theta (1- \theta) )/(n+2)^2

For low values of n and large values of \theta, the risk for \hat{\theta} will be larger than that of \tilde{\theta}, but for high values of n, the risk for \hat{\theta} will be lesser than that of \tilde{\theta}.